A 3-kg block moving to the right on a frictionless table at 4 m/s makes a head on collision with another 3-kg block moving 5 m/s to the left.


*If the collision is completely inelastic, find the final velocity of the blocks

6.


*If the collision is completely elastic, find the final velocities of the blocks.Required to answer. Single choice.


*If half the initial energy is dissipated in the collision, find the final velocities of the blocks.Required to answer. Single choice.








1
Expert's answer
2021-02-05T11:47:55-0500

A. If the collision is completely inelastic, find the final velocity of the blocks using momentum conservation:


"p_i=p_f,\\
mv-Mu=(m+M)s, \\\space\\
s=\frac{mv-Mu}{m+M}=\frac{3\cdot4-3\cdot5}{3+3}=-0.5\text{ m/s}."

The two blocks will move to the left at 0.5 m/s as one.


B. If the collision is completely elastic, find the final velocities of the blocks.

In this scenario, we will also introduce conservation of energy:


"p_i=p_f,\\
mv_1-Mu_1=mv_2+Mu_2, \\\space\\
\frac12mv_1^2+\frac12Mu_1^2=\frac12mv_2^2+\frac12Mu_2^2."

In this system of equations, we have two undefined values, namely, "v_2, u_2." The solution is


"v_2=-5\text{ m/s},\\
u_2=4\text{ m/s}."

C. If half the initial energy is dissipated in the collision, find the final velocities of the blocks. Use the reasonings from part B, but include the lost energy condition: just 50% of the energy goes for final kinetic energy of motion.


"p_i=p_f,\\
mv_1-Mu_1=mv_2+Mu_2, \\\space\\
\frac12\bigg(\frac12mv_1^2+\frac12Mu_1^2\bigg)=\frac12mv_2^2+\frac12Mu_2^2."

The solution:


"v_2=-3.66\text{ m/s}, \\
u_2=2.66\text{ m/s}."